A 1.27g gas occupies a volume of 1.13L at 25°C and 746 mm Hg. The gas may be
(1) CO
(2) CO2
(3) CH4
(4) N2O

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
PV = n RT
 
PV = (w /m) RT
 
(746/760) * 1.13 = (1.27 / m) 0.0821 * 298
 
m = 1.27 * 0.0821 * 298 *760/ (1.13 * 746)
 
m = 28.012
 
Hence gas is CO